OpenStudy (anonymous):

PLEASE PLEASE HELP Two quadratic functions are shown. Function 1: f(x) = 3x2 + 6x + 7 Function 2: x g(x) −2 13 −1 7 0 3 1 7 Which function has the least minimum value and what are its coordinates?

OpenStudy (rajat97):

you can find the minimum value of a quadratic function by using the formula -b/2a if we assume that the equation is given in the form y=ax^2 + bx +c then, try to plot the values of the second function roughly on a graph paper or even a simple paper will do good with some estimation and then, you can get the answer

OpenStudy (anonymous):

I dont know how to use that formula. Im new to Algebra.

OpenStudy (rajat97):

okay! no problem assume that we have a quadratic equation of the form y=ax^2 + bx + c now, compare the equation y=ax^2 + bx + c (we call it the standard equation) to the given equation, i.e. f(x) = 3x2 + 6x + 7 so we get a=3, b=6, c=7 now put the values of a abd b in the formula that i have specified in the above post if you want it's derivation, i can help you with it

OpenStudy (anonymous):

y=3x^2+6x+7?

OpenStudy (rajat97):

yes

OpenStudy (lynfran):

i think this question here just require you to input "x" they already give they g(x) solution for various values of "x" so; when x=-2 f(-2)=3(-2)^2+6(-2)+7..=7 .and ....g(-2)=13 when x=-1 f(-1)=3(-1)^2+6(-1)+7....=4 and ...g(-1)=7 when x=0 f(0)=3(0)^2+6(0)+7..=7 ...and g(0)=3 when x=1 f(1)=3(1)^2+6(1)+7 ...= 16 and g(1)=7 so we compare to see which function has the least minimum value..

OpenStudy (anonymous):

so function 1 has the least minimum value and its coordinates are 0,7 @LynFran

OpenStudy (lynfran):

no that would be function 2... with coordinate (0,3)

OpenStudy (anonymous):

wow thankyou so much! @LynFran

OpenStudy (anonymous):

Can u help me with more?

OpenStudy (lynfran):

welcome

OpenStudy (lynfran):

ok

OpenStudy (anonymous):

2)A function is shown in the table. x g(x) −2 2 −1 −3 0 2 1 17 Which of the following is a true statement for this function? A: The function is increasing from x=-2 to x=-1 B: The function is increasing from x=0 to x=1 C: The function is decreasing from x=-1 to x=0 D: The function is decreasing from x=0 to x=1

OpenStudy (anonymous):

@LynFran

OpenStudy (lynfran):

well A, C and D are incorrect so...

OpenStudy (anonymous):

lol guess its b

OpenStudy (anonymous):

thankyou again so much.

OpenStudy (lynfran):

welcome

OpenStudy (rajat97):

thank you @UnkleRhaukus for the medal